# A Bayesian approach with Gaussian priors to the inverse problem of source identification in elliptic PDEs

### v2

@@ 1 Introduction
Observation schemes as in equation (1) are found in a variety of scientific fields and engineering applications, including medical imaging [9], geophysics [39], acoustics [12] and finance [7]. For example, Computerised Tomography (CT), a technique to obtain detailed images of the human body and information about the density variation of the tissues, is based on a mathematical model (related to the ‘Radon transform’) for the absorption of $X$-rays. Similar concepts underpin many other medical imagining techniques, such as Magnetic Resonance (MR) and Positron Emission Tomography (PET); see [9] for further references.
@@
The Bayesian approach to inverse problems dates back at least to the 1980’s, with early seminal work laid out in [42, 30, 31] among the others, and has since gained enormous popularity across several applied fields. See the monographs [24, 41], as well as the more recent reviews [8, 6], where further references can be found. Over the last decade, a large number of articles have investigated the theoretical recovery performance of Gaussian priors in linear inverse problems. We refer in particular to [28, 29, 4, 26], and also mention [1, 35, 19, 34, 17] for results in nonlinear problems. Recently, Giordano and Kekkonen [18], building on earlier findings by [33], identified general conditions on the forward operator, on the ground truth and on the prior distribution under which ‘semiparametric’ Bernstein-von Mises theorems can be obtained, characterising the asymptotic shape of the posterior distribution of a large collection of linear functionals of the unknown. For the problem of estimating the source $f$ in the elliptic PDE (2) from observations $Y^{\varepsilon}$ arising as in (1), they showed that a wide class of ‘standard’ centred Gaussian process priors (such as the one associated to the commonly used Matérn kernel) yield, in the small noise limit $\varepsilon\to 0$ and under the frequentist assumption that the data have been generated by a fixed ground truth $f_{0}$, valid and optimal estimation and uncertainty quantification, via the posterior mean estimate and credible intervals centred around it, for one-dimensional aspects of the unknown source function.
@@ 2.1 Details on the Bayesian model
Throughout, let $\mathcal{O}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, be a non-empty, open and bounded set with smooth boundary $\partial\mathcal{O}$. For a fixed, smooth and positive ‘diffusion coefficient’ $c\in C^{\infty}(\overline{\mathcal{O}})$, $\inf_{x\in\mathcal{O}}c(x)>0$, let $G:L^{2}(\mathcal{O})\to L^{2}(\mathcal{O})$ be the (linear) source-to-solution map associated to the elliptic PDE (2); see Appendix B in [18] for details. For such $G$, and a given noise level $\varepsilon>0$, consider observations $Y^{\varepsilon}$ arising as in (1) for some unknown $f\in L^{2}(\mathcal{O})$, where $W$ is a white noise process indexed by $L^{2}(\mathcal{O})$ defined on some probability space $(\Omega,\mathcal{S},\Pr)$, that is the centred Gaussian process $(W(g),\ g\in L^{2}(\mathcal{O}))$ with covariance $\textnormal{E}[W(g_{1})W(g_{2})]=\langle g_{1},g_{2}\rangle_{2}$. Throughout most of the article, we will regard $\varepsilon$ as known. In practice, it can often be replaced by an estimate (cf. Section 3.3.2). As described, for example, in Chapter 1 of [16], observing $Y^{\varepsilon}$ is understood as observing a realisation of the Gaussian process $(Y^{\varepsilon}(g),\ g\in L^{2}(\mathcal{O}))$ with mean $\textnormal{E}[Y^{\varepsilon}(g)]=\langle f,g\rangle_{2}$ and covariance $\mathbbm{C}\textnormal{ov}[Y^{\varepsilon}(g_{1}),Y^{\varepsilon}(g_{2})]=\langle g_{1},g_{2}\rangle_{2}$. Such observation scheme serves as a convenient continuous counterpart of the inverse regression model
@@
2.2 Theoretical guarantees on estimation and uncertainty quantification
@@
where $R_{\varepsilon,a}>0$ is the $(1-a/2)\%$-quantile of the one-dimensional posterior distribution of $\langle f,\psi\rangle_{2}|Y^{\varepsilon}$, so that
@@ 2.3 Examples of Gaussian priors
then satisfies $\mathbb{H}^{0}=L^{2}(\mathcal{O})$ (with equality of norms) and the continuous (strict) embedding $\mathbb{H}^{\alpha}\subset H^{\alpha}(\mathcal{O})$ for all $\alpha>0$ [43, p. 472]. In fact, it holds that $\|f\|_{\mathbb{H}^{\alpha}}\simeq\|f\|_{H^{\alpha}}$ for all $f\in\mathbb{H}^{\alpha}$ and $\alpha\geq 0$.
@@
corresponding to the Karhunen-Loève expansions of certain commonly used Gaussian process priors with covariance kernel given by an inverse power of the Laplacian [40, Section 2.4]. By Weyl’s asymptotics, we have
@@
A second example of interest involves stationary Gaussian processes defined via a covariance kernel of choice. In particular, the Matérn kernel is widely used in applications [36, Section 4.2].
@@ 3.2 Posterior inference with the Matérn process prior
The same conjugate computation as the one outlined in Section 3.1.1 can then be carried out, leading to the Gaussian posterior distribution $\mathbf{f}|\mathbf{Y}\sim N(\bar{\mathbf{f}}_{n},\mathbf{C}_{n})$, with posterior mean and covariance matrix respectively given by
@@ 3.3.2 Inference with unknown noise variance
Table 5: Inferential results for the difference-based estimation $\hat{\sigma}_{n}$ of the noise standard deviation and the the empirical Bayes posterior mean estimator $\hat{f}_{n}$, for increasing sample size.

### v3

@@ 1 Introduction
Observation schemes as in equation (1) are found in a variety of scientific fields and engineering applications, including medical imaging [9], geophysics [40], acoustics [12] and finance [7]. For example, Computerised Tomography (CT), a technique to obtain detailed images of the human body and information about the density variation of the tissues, is based on a mathematical model (related to the ‘Radon transform’) for the absorption of $X$-rays. Similar concepts underpin many other medical imagining techniques, such as Magnetic Resonance (MR) and Positron Emission Tomography (PET); see [9] for further details and references.
@@
The Bayesian approach to inverse problems dates back at least to the 1980’s, with early seminal work laid out in [43, 31, 32] among the others, and has since gained enormous popularity across several applied fields. See the monographs [25, 42], as well as the more recent reviews [8, 6], where further references can be found. Over the last decade, a large number of articles have investigated the theoretical recovery performance of Gaussian priors in linear inverse problems. We refer in particular to [29, 30, 4, 27, 20], and also mention [1, 36, 19, 35, 17] for results in nonlinear problems. Recently, Giordano and Kekkonen [18], building on earlier findings by [34], identified general conditions on the forward operator, on the ground truth and on the prior distribution under which ‘semiparametric’ Bernstein-von Mises theorems can be obtained, characterising the asymptotic shape of the posterior distribution of a large collection of linear functionals of the unknown. For the problem of estimating the source $f$ in the elliptic PDE (2) from observations $Y^{\varepsilon}$ arising as in (1), they showed that a wide class of ‘standard’ centred Gaussian process priors (such as the one associated to the commonly used Matérn kernel) yield, in the small noise limit $\varepsilon\to 0$ and under the frequentist assumption that the data have been generated by a fixed ground truth $f_{0}$, valid and optimal estimation and uncertainty quantification, via the posterior mean estimate and credible intervals centred around it, for one-dimensional aspects of the unknown source function.
@@ 2.1 Details on the Bayesian model
Throughout, let $\mathcal{O}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, be a non-empty, open and bounded set with smooth boundary $\partial\mathcal{O}$. For a fixed, smooth and positive ‘diffusion coefficient’ $c\in C^{\infty}(\overline{\mathcal{O}})$, $\inf_{x\in\mathcal{O}}c(x)>0$, let $G:L^{2}(\mathcal{O})\to L^{2}(\mathcal{O})$ be the (linear) source-to-solution map associated to the elliptic PDE (2); see Appendix B in [18] for details. For such $G$, and a given noise level $\varepsilon>0$, consider observations $Y^{\varepsilon}$ arising as in (1) for some unknown $f\in L^{2}(\mathcal{O})$, where $W$ is a white noise process indexed by $L^{2}(\mathcal{O})$ defined on some probability space $(\Omega,\mathcal{S},\Pr)$, that is the centred Gaussian process $(W(g),\ g\in L^{2}(\mathcal{O}))$ with covariance $\textnormal{E}[W(g_{1})W(g_{2})]=\langle g_{1},g_{2}\rangle_{2}$. Throughout most of the article, we will regard $\varepsilon$ as known. In practice, it can often be replaced by an estimate (cf. Section 3.3.2). As described, for example, in Chapter 1 of [16], observing $Y^{\varepsilon}$ is understood as observing a realisation of the Gaussian process $(Y^{\varepsilon}(g),\ g\in L^{2}(\mathcal{O}))$ with mean $\textnormal{E}[Y^{\varepsilon}(g)]=\langle G(f),g\rangle_{2}$ and covariance $\mathbbm{C}\textnormal{ov}[Y^{\varepsilon}(g_{1}),Y^{\varepsilon}(g_{2})]=\langle g_{1},g_{2}\rangle_{2}$. Such observation scheme serves as a convenient continuous counterpart of the inverse regression model
@@
2.2 Theoretical guarantees for estimation and uncertainty quantification
@@
where $R_{\varepsilon,a}>0$ is the $(1-a/2)\%$-quantile of the one-dimensional (Gaussian) posterior distribution of $\langle f,\psi\rangle_{2}|Y^{\varepsilon}$, so that
@@ 2.3 Examples of Gaussian priors
then satisfies $\mathbb{H}^{0}=L^{2}(\mathcal{O})$ (with equality of norms) and the continuous (strict) embedding $\mathbb{H}^{\alpha}\subset H^{\alpha}(\mathcal{O})$ for all $\alpha>0$ [44, p. 472]. In fact, it holds that $\|f\|_{\mathbb{H}^{\alpha}}\simeq\|f\|_{H^{\alpha}}$ for all $f\in\mathbb{H}^{\alpha}$ and $\alpha\geq 0$.
@@
corresponding to the Karhunen-Loève expansions of certain commonly used Gaussian process priors with covariance kernel given by an inverse power of the Laplacian [41, Section 2.4]. By Weyl’s asymptotics, we have
@@
A second example of interest involves stationary Gaussian processes defined via a covariance kernel of choice. In particular, the Matérn kernel is widely used in applications [37, Section 4.2].
@@ 3.2 Posterior inference with the Matérn process prior
The same conjugate computation as the one outlined in Section 3.1.1 can then be carried out, leading to the Gaussian posterior distribution $\mathbf{f}|\mathbf{Y}\sim N_{M}(\bar{\mathbf{f}}_{n},\mathbf{C}_{n})$, with posterior mean and covariance matrix respectively given by
@@ 3.3.2 Inference with unknown noise variance
Table 5: Inferential results for the difference-based estimator $\hat{\sigma}_{n}$ of the noise standard deviation and the the empirical Bayes posterior mean estimator $\hat{f}_{n}$, for increasing sample sizes.